# Evidence of Scaling Regimes in the Hopfield Dynamics of Whole Brain Model

### v1

@@ IV.1 Scaling exponents as a function of the decay length
Several comments are in order. First, we observe a clear increasing trend of $\alpha$ with $\delta$, meaning that small decay lengths, i.e. localised connectivity, promotes irregular patterns. For instance, below $\delta\sim 4$ mm, the scaling exponent is basically zero, corresponding to complete randomness. This was checked independently, by running simulations in which the couplings $J_{ij}$ were randomized though keeping the same distribution of the elements of the matrix $J$ (flat orange line at the bottom). We were able to simultaneously randomize the couplings $J_{ij}$, and keep the same value distribution by randomly shuffling the $J_{ij}$ values. At the opposite end, with $\delta=10$ mm, we measure $\alpha\sim 1.4$, which is still well below the smoothness threshold $\alpha_{S}=2$, but above the random walk value $\alpha_{R}=1$, and much above the Kolmogorov turbulence value $\alpha_{T}=2/3$, which is in turn larger than $\alpha_{D}\sim 1/2$, the value obtained by Deco et al. at $\delta\sim 5.55$ mm. Importantly, in the intermediate regime, around the physiological value $\delta=5.55$ mm, the present Hopfield model delivers $\alpha_{H}\sim 2/5$, different but still close to Deco et al. value.
@@
Figure 4 shows how the pair correlation for different values of $\delta$. For $\delta<5$ mm, the pair correlation shows a basically uncorrelated activity, except at short range. For $\delta=5.26$ mm, long-range pair correlations start to emerge and finally, for larger $\delta$, the long-range pair correlations become even stronger. In the limit of very large $\delta$, stationary states emerge in which the system freezes into a single fully-ordered state, with all nodes states either $+1$ or $-1$.
@@
Our analysis confirms that there is indeed a scaling regime in the organization of the brain patterns which reminds fluid turbulence, yet with a higher degree of spikiness. Indeed, by sticking to the physiological value $\delta=5.55$ mm, the Hopfield brain appears more irregular (lower scaling exponent) than a random walk and also of homogeneous incompressible turbulence. The most important feature, though, is the sigmoidal dependence of the relation $\alpha=\alpha(\delta)$ with the rampup region between about $4$ and $7$ mm, centered around $\delta\sim 5.9$. This is indeed a physiologically relevant scale of local regions in the brain. Furthermore, this rampup region unveils the major sensitivity of the scaling exponent to small changes of the decay length, indicating that small changes of the decay length lead to fairly different scaling regimes. In particular, for $\delta$ below $4$ mm, the activity appears to be nearly chaotic, whereas for $\delta$ above $8$ mm, it becomes kind of “rigid”, i.e. globally correlated.
A word of caution must be spent on the statistical accuracy of the scaling exponents, which for the present case of $40$ realizations, is around $25$ percent.
Figure 4: Pair-correlation functions $B(d)$ for different $\delta$s. At small $\delta$ the signal is largely uncorrelated, while at increasing $\delta$, correlations start to emerge, although affected by significant amount of statistical fluctuations.
V Thresholding the connections
Figure 5: The average slope of $\log S(d)$ for different dilution levels $\alpha(\rho)$, normalized to its threshold-free value $\alpha(\rho)$ at that given $\delta$. The dilution is measured as the fraction of disconected pairs $i,j$, $J_{ij}$ elements with value 0. The second $y$-axis on top shows the cutoff distance in mm for wich at a certain dilution $\rho$ the pairs at a larger or euql distance are disconected. The differently colored curves indicate simulations with different $\delta$ values. The graph shows that within the measurement error, the system is not affected by the dilution of the edges, up to the removal of more than the $95\%$ of the connections, in descending order of distance. This shows that, in view of the exponential connectivity $J_{ij}=e^{-d_{ij}/\delta}$, the onset of collective patterns is sustained mostly by the close connections, within about $4\delta$.
@@ V Thresholding the connections
Having assessed that delocalization (large $\delta$s) promotes a correlated response, it is natural to inquire the role of short versus long-range couplings in promoting the correlated patterns sustaining the scaling regime.
@@
To this purpose, we performed a series of simulations by progressively removing pairs of nodes whose coupling strengths lie below a given threshold, namely $J_{ij}<J_{th}$, where the coupling strengths vary in the range $[0,1]$. Since $J_{ij}$ decays exponentially with the pair distance, it is clear that the running threshold removes first the distant nodes and then the less distant ones, upon increasing the threshold, thereby implementing a “correlated dilution” process. More precisely, a threshold $J_{th}$ excludes pairs beyond a distance
@@
thereby retaining only a fraction $1-J_{th}$ of the full set of interacting nodes. The result of this progressive dilution on the scaling exponent are reported in Figure 5. Where the dilution $\rho$ is defined as the fraction of disconnected nodes $i,j$, $J_{ij}=0$, and the resolt on the scaling exponent is masured comparing the undiluted exponent $\alpha(0)$ with the exponent $\alpha(\rho)$ at that dilution. Since the dilution is obtaned thresholding in order the $J_{ij}$ values untill we get a certain fraction of disconnected nodes, and since eq. 12 we get that to a certain dilution corresponds a certain cutoff distance above which all couples are disconnected.
@@
From this figure, a smooth trend at increasing $\rho$ is observed, with the scaling exponent starting to display significant decrease above $\rho=\left[0.95,0.98\right]$, when just 2 or 5 percent of the nodes are left, corresponding to $d_{ij}>4.6\delta$. With $J_{th}=0.10$, corresponding to $d_{ij}>-\delta ln(0.1)\sim 0.46\delta$, the scaling exponent is basically halved, indicating a major loss of correlation. This is expected since $J_{th}=0.10$ cuts out all but shortest range interactions.
The conclusion of this dilution analysis is that the brain response is carried almost entirely by interactions up to about five decay lengths, namely about $1/5$ of the size of the global brain. This sounds reasonable, as it strikes a plausible compromise between short-range ($d<4\delta$) and long-range ($d>6\delta$) interactions. It is interesting to observe that a “solid” (strongly correlated) brain, such as the one that results for $\delta>10$ mm, would be too homogeneous to be able to store and process the amount of information required to function properly.
On the opposite side, for too short connectivity length, the brain would behave like a “gas” of disconnected neurons, hence incapable of collective behavior, which is key to its proper functioning
@@
It appears like the physiological length is achieving an optimal compromise between these two opposite, order-disorder, trends. With a daring but conducive metaphor, we could speculate that the brain works in a sort of turbulent liquid-like state, although more spiky than actual turbulent fluids.
VI Conclusions
@@
It is shown that an Hopfield recurrent neural network, informed by experimentally derived brain topology, recovers the scaling picture recently proposed by Deco Deco2020 , according to which the process of information exchange within the human brain shows spatially correlated patterns qualitatively similar to those displayed by turbulent flows.
Our analysis confirms the initial finding by Deco2020 in near-quantitative form, predicting very similar scaling exponent for the same value of the connectivity decay length. This is pretty remarkable, given the very different nature of the two mathematical models and provides a strong hint at the robustness of the scaling picture.
@@
It is further observed that the scaling exponents are smaller than for the case of turbulence, indicating that the collective activity of the brain is more irregular (spiky) than for a homogeneous incompressible turbulent fluid.
Importantly, the scaling exponents show a steep dependence on the spatial range of neural connections, hence the specific value of the decay length has a major impact on the degree of irregularity (spikiness) of the collective patterns.
@@
This provides a strong indication that there is an optimal connectivity length at which the brain shows non-smooth functional behavior. This appears to happen at a scale around $5,5$ mm, hence well below the global size of the brain Longer/shorter connectivities would lead to excessive smoothness/roughness respectively. Interestingly, a dilution analysis shows that the Hopfield model brain remains functional by removing links above about six decay lengths, namely $25-30$ mm, corresponding to about one fifth of the size of the global brain. This indicates that, in terms of connectivity decay length, the brain seems to function in a sort of intermediate ”turbulent liquid-like” state , whose essential connections lie in near geometrical mean between the decay length and the global size of the brain. This turbulent liquid-like state appears to be more spiky than ordinary homogeneous incompressible turbulent fluids.
VII Acknowledgments
@@
One of the authors (SS) wishes to acknowledge financial support from the Bertarelli Foundation during his stay at the Neurobiology Department of the Harvard Medical School. The authors wish to acknowledge illuminating discussions with John A. Assad, Haim Sompolinsky, and Bernardo Sabattini. GG contributes to code writing and simulations. GG, SS, and GR conceived the Idea, designed the experiments, interpreted the data, and wrote the manuscript.

### v2

@@ Introduction
Several comments are in order. First, we observe a clear increasing trend of $\alpha$ with $\delta$, meaning that small decay lengths, i.e. localised connectivity, promotes irregular patterns. For instance, below $\delta\sim 4$ mm, the scaling exponent is basically zero, corresponding to complete randomness. This was checked independently, by running simulations in which the couplings $J_{ij}$ were randomized though keeping the same distribution of the elements of the matrix $J$ (flat orange line at the bottom). We were able to simultaneously randomize the couplings $J_{ij}$, and keep the same value distribution by randomly shuffling the $J_{ij}$ values. At the opposite end, with $\delta$=10 mm, we measure $\alpha$$\sim$1.2, which is still well below the smoothness threshold $\alpha_{S}$=2, but above the random walk value $\alpha_{R}$=1, and much above the Kolmogorov turbulence value $\alpha_{T}$=2/3, which is in turn larger than $\alpha_{D}$$\sim$1/2, the value obtained by Deco et al.  (?) at $\delta$$\sim$5.55 mm. Importantly, in the intermediate regime, around the physiological value $\delta$=5.55 mm, the present Hopfield model delivers $\alpha_{H}$$\sim$2/5, different but still close to Deco et al. value. A visual extrapolation of Fig. 2A indicates that the scaling regime would remain non-smooth ($\alpha<2$) even in the limit of an ”infinite-brain”, $\delta\to\infty$), the global brain ($d\sim 25\delta$) being pretty close to the infinite-brain limit. Note however that for $\delta>10$ mm we start getting cases in which $B(d)\sim B(0)$, so that the numerical calculation of the exponent diverges and becomes less and less accurate (a fully correlated signal with $B(d)=B(0)$ yields formally $\alpha\to\infty$.)
@@
Figure 3 shows the pair correlation for different values of $\delta$. For $\delta<5$ mm, the pair correlation shows a basically uncorrelated activity, except at short range. For $\delta=5.26$ mm, long-range pair correlations start to emerge and finally, for larger $\delta$, the long-range pair correlations become even stronger. In the limit of very large $\delta$, stationary states emerge in which the system freezes into a single fully-ordered state, with all nodes states either $+1$ or $-1$. It is important to consider that in the fully-ordered state with no noise $\alpha$ is infinite.
@@
Our analysis confirms that there is indeed a intermediate scaling regime in the organization of the brain patterns which reminds fluid turbulence, yet with a higher degree of spikiness. Indeed, by sticking to the physiological value $\delta=5.55$ mm, the Hopfield brain appears more irregular (lower scaling exponent) than a random walk and also of homogeneous incompressible turbulence. The most important feature, though, is the sigmoidal dependence of the relation $\alpha(\delta)$ with the rampup region between about $4$ and $7$ mm, centered around $\delta\sim 5.9$. This is indeed a physiologically relevant scale of local regions in the brain. Furthermore, this rampup region unveils the major sensitivity of the scaling exponent to small changes of the decay length, indicating that small changes of the decay length lead to fairly different scaling regimes. In particular, for $\delta$ below $4$ mm, the activity appears to be nearly chaotic, whereas, for $\delta$ above $8$ mm, it becomes kind of “rigid”, i.e. globally correlated.
Thresholding the connections
@@
Having assessed that delocalization (large $\delta$s) promotes a correlated response, it is natural to inquire about the role of short versus long-range couplings in promoting the correlated patterns sustaining the scaling regime.
@@
To this purpose, we performed a series of simulations by progressively removing pairs of nodes whose coupling strengths lie below a given threshold, namely $J_{ij}<J_{th}$, where the coupling strengths vary in the range $[0,1]$. Since $J_{ij}$ decays exponentially with the pair distance a threshold $J_{th}$ excludes pairs beyond a distance
@@
thereby retaining only a fraction $1-J_{th}$ of the full set of interacting nodes. Thus, to a certain dilution corresponds a certain cutoff distance above which all couples are disconnected. The result of this progressive dilution on the scaling exponent are reported in Figure 4. Here the dilution $\rho$ is defined as the fraction of disconnected nodes $i,j$, $J_{ij}=0$, and the result on the scaling exponent is measured comparing the undiluted exponent $\alpha(0)$ with the exponent $\alpha(\rho)$ at that dilution.
@@
From this figure, a smooth trend at increasing $\rho$ is observed, with the scaling exponent starting to display significant decrease above $\rho=\left[0.95,0.98\right]$, when just 2 or 5 percent of the nodes are left, corresponding to $d_{ij}>4.6\delta$. With $J_{th}=0.10$, corresponding to $d_{ij}>-\delta ln(0.1)\sim 0.46\delta$, the scaling exponent is basically halved, indicating a major loss of correlation. This is expected since $J_{th}=0.10$ cuts out all but shortest range interactions. The conclusion of this dilution analysis is that the brain response is carried almost entirely by interactions up to about five decay lengths, namely about $1/6$ of the size of the global brain. This sounds reasonable, as it strikes a plausible compromise between short-range ($d<4\delta$) and long-range ($d>6\delta$) interactions.
It is interesting to observe that a “solid” (strongly correlated) brain, such as the one that results for $\delta>10$ mm, and low dilution $\delta$, would be too homogeneous to be able to store and process the amount of information required to function properly. On the opposite side, for too short connectivity length, the brain would behave like a “gas” of disconnected nodes, hence incapable of collective behavior, which is key to its proper functioning. It appears like the physiological length is achieving an optimal compromise between these two opposite, order-disorder, trends.
@@
With a daring but conducive metaphor, we could speculate that the brain works in a sort of turbulent liquid-like state, although more spiky than actual turbulent fluids.
Changing The Number of Parcels
To assess the impact of the parcellation size on the scaling picture discussed in this paper, we gathered Schaefer 2018 Parcellations with different sizes $N$ and we run numerical simulations with $N_{r}=1000$ replicas and different values of $\delta$. This allowed us to inspect the relation between $\alpha$ and $\delta$ for different values of the size $N$. The result is the relation reported in Fig. 5, which shows a progressive regression of the transition regions towards smaller values of $\delta$, combined with a corresponding steepening of the transition.
To provide a quantitative measure of this process, we fitted the points at given values of $N$ with the sigmoid function:
$\alpha(\delta)=\frac{\alpha_{\infty}}{1+\exp{(-k(\delta-\delta_{o}))}}.$ (7)
where $\alpha_{\infty}$ represents the plateau value of the scaling exponent at infinite decay length, $\delta_{0}$ is the transition scale and $k$ is the steepness of the transition region.
The sigmoid function captures the functional relation of the points reasonably well even though, as we discussed earlier on, the values $\alpha(\delta)$ do not fully converge in the top branch because for $\delta>10$ mm fully aligned configurations prevail, with correspondingly divergent values of $\alpha$.
Thus, we fitted the function in two steps: first, we estimated $\alpha_{\infty}$ on the plateau points for $N=800,900,1000$. Then, we obtained $\delta_{0}$ and $k$ by fitting the sigmoid function $\alpha(\delta)$ on all points using a nonlinear least squares fit. We finally obtained $\delta_{0}$ and $k$ as a function of $N$ via a linear least square fit in loglog scale, to deliver $\delta_{0}\sim N^{-0.379}$ and $k\sim N^{0.328}$, with a $R^{2}$-value $0.999$ and $0.925$, respectively
Based on the above scaling of the parameters, one would conclude that the sigmoid tends to a sign function centered in zero in the thermodynamic limit $N\to\infty$. In plain words, the transition length goes to zero and the transition region becomes infinitely steep. Hence, in the infinite $N$ thermodynamic limit, the system is fully correlated (“rigid”) even for vanishing decay lengths, an extreme instance of long-range ordering.
But the brain is not infinite, hence it is plausible to expect that there should exist optimal parcellation size $N_{opt}$ providing a correspondingly optimal Hopfield mass model approximation to the empirical correlations provided by a realistic model of structural connectivity. Due to its inherently coarse-grained nature, above $N_{opt}$, the Hopfield network whole brain mass model is no longer expected to provide a trustworthy approximation of the collective brain dynamics.
Even if we cannot pin down such an optimal value, we expect it to lie between $N=10^{3}$, the largest parcellation presently available to us and $N<10^{6}$, which provides a plausible upper bound to the number of functioning ”modules” in the human brain (number of neurons, $10^{11}$, divided by the typical module size, $10^{5}$ neurons).
Given the actual scalings provided by our fit, a three orders of magnitude increase in the parcellation size would lead to about ten times smaller values of both $\delta_{0}$ and thickness $1/k$, thus placing the transition length $\delta_{0}$ at about $0.5$ mm, with a strong sensitivity of the scaling exponent confined within a region of lengthscale about $0.1$ mm wide. In other words, going to larger parcellations makes the scaling exponent increasingly sensitive to small changes of the decay length.
Future experimental developments will allow us to explore finer parcellations with larger $N$ values and test the predictions made in this work. Furthermore, from these and other independent considerations  (?), it is plausible to speculate that for $N>10^{3}$, the connectivity may no longer be approximated by a decaying exponential.
All of the above, point to the importance of obtaining larger parcellations, providing further constraints to improve future mesoscale models of the global brain activity.
Conclusions
@@
In conclusion, we have shown that an Hopfield recurrent neural network, informed by experimentally derived brain topology, recovers the scaling picture recently proposed by Deco  (?), according to which the process of information exchange within the human brain shows spatially correlated patterns qualitatively similar to those displayed by turbulent flows. Our analysis confirms the initial finding by  (?) in near-quantitative form, predicting a very similar scaling exponent ($2/5$ instead of $1/2$) for the same value of the connectivity decay length. Given the very different nature of the two mathematical models, this provides a significant hint at the robustness of the scaling picture.
@@
It is further observed that the scaling exponents are smaller than for the case of turbulence, indicating that the collective activity of the brain is more irregular (spikier) than homogeneous incompressible turbulence.
The scaling exponents show a steep dependence on the spatial range of neural connections, hence the specific value of the decay length has a major impact on the degree of irregularity (spikiness) of the collective patterns. Such sensitivity is also predicted to increase by increasing the size of the parcellation. This provides a strong indication towards the existence of an optimal connectivity length $\delta_{0}$ at which the brain shows non-smooth functional behavior, meaning that longer/shorter connectivity lengthscales would lead to excessive smoothness/roughness respectively. With $N=10^{3}$ parcels, this lengthscale is about $\delta_{0}\sim 5.5$ mm, well below the global size of the brain. For larger parcellations, both the transition length $\delta_{0}$ and the width of the transition region are predicted to decrease approximately like $N^{-1/3}$, which highlights the importance of assessing the optimal parcellation size.
@@
Finally, a dilution analysis shows that, again with $N=10^{3}$, the Hopfield model brain remains functional after removing links above about six decay lengths, namely $25-30$ mm, corresponding to about one-fifth of the size of the global brain. This indicates that, in terms of connectivity decay length, the brain seems to function in a sort of intermediate “turbulent liquid-like” state, whose essential connections lie in near geometrical mean between the decay length and the global size of the brain.
The Hopfield network model and the webXR parcels activity 3d representation will be available to access in open cloud servers ( e.g. H2IOSC data centers).
Acknowledgments
@@
One of the authors (SS) wishes to acknowledge financial support from the Bertarelli Foundation during his stay at the Neurobiology Department of the Harvard Medical School. The authors wish to acknowledge illuminating discussions with John A. Assad, Haim Sompolinsky, and Bernardo Sabattini.
GG wishes to acknowledge financial support H2IOSC Project - Humanities and cultural Heritage Italian Open Science Cloud funded by the European Union – NextGenerationEU – NRRP M4C2 - Project code IR0000029 - CUP B63C22000730005.
Supplementary materials
Supplementary Text composed of the three paragraphs: i) the turbulent-like whole-brain model, ii) Amplitude of turbulence and scaling regime, iii) Structural connectivity.
Figure S1
Fig. 1. Structure factor $S_{2}(d)$ for $N=1000$ nodes and $N_{r}=1000$ realizations respectively at $\delta=5.55$ mm (A), and $\delta=5.88$ mm (B). The values $S_{2}(d)$ were binned over equal sized intervals. The dashed lines indicate different $\alpha$ values: $\alpha=2/3$ which corresponds to turbulence, $\alpha=1/2$, the value obtained in  (?), and $\alpha=2/5$. We estimate the scaling exponent $\alpha$ with a linear regression of $\log(S(d))\sim\alpha\log(d)+\beta$ in the interval range $[2.7-33.1]$ mm, as used in  (?), and we obtain $\alpha\approx 2/5$ for $\delta=5.55$ mm, and $\alpha\approx 2/3$ for $\delta=5.88$ mm.
Fig. 2. A) The blue line shows the average slope of $\alpha=\log S(d)/\log d$ for different $\delta$s, given the connectivity matrix $J_{ij}=\exp{(-d_{i,j}/\delta)}$. The horizontal dotted lines correspond to $\alpha_{RW}=1$ and $\alpha_{T}=2/3$, corresponding to random walk and homogeneous incompressible turbulence, respectively, while the dot-dashed horizontal line corresponds to Deco et al. “brain” value $\alpha_{D}=1/2$. The orange line shows the function $\alpha(\delta)$ with a connectivity matrix obtained by randomly shuffling all the weights $J_{ij}$, which results into a basically uncorrelated signal, indicating the importance of the metric structure of the weights. The plot stops at $\delta=10$ mm, a region where the signal is highly correlated ($\alpha\sim 1.2$), but still below the smooth regime marked by $\alpha=2$. Conversely, a brain below $4$ mm, would exhibit “chaotic” behaviour, with scaling exponents close to zero, i.e. no scaling regime. The bottom panels show the neuron activity mapped on the Schaefer’s cerebral cortical parcellation atlas coordinates for (B) $\delta=4$ (“chaotic”), (C) $\delta=5.99$, and (D) $\delta=6.66$ (random walk), respectively.
Fig. 3. Pair-correlation functions $B(d)$ for different $\delta$s. At small $\delta$ the signal is largely uncorrelated, while at increasing $\delta$, correlations start to emerge, although affected by significant amount of statistical fluctuations.
Fig. 4. The average slope of $\log S(d)$ for different dilution levels $\alpha(\rho)$, normalized to its threshold-free value $\alpha(0)$ at that given $\delta$. The dilution is measured as the fraction of disconnected pairs $i,j$, $J_{ij}$ elements with value 0. The second $y$-axis on top shows the cutoff distance in mm for which at a certain dilution $\rho$ the pairs at a larger or equal distance are disconnected. The differently colored curves indicate simulations with different $\delta$ values. The graph shows that within the measurement error, the system is not affected by the dilution of the edges, up to the removal of more than the $95\%$ of the connections, in descending order of distance. This shows that, in view of the exponential connectivity $J_{ij}=e^{-d_{ij}/\delta}$, the onset of collective patterns is sustained mostly by the close connections, within about $4\delta$.
Fig. 5. The average slope of $\alpha$ as a function of $\delta$ for Schaefer 2018 Parcellations with different sizes $N$. Given each size $N$, we fitted the points with the sigmoid functions in Eq. (7).